Question 2 The built-up shaft consists of a pipe AB and solid rod BC. The pipe has an inner diameter of 20 mm and outer diameter of 28 mm. The rod has a diameter of 12 mm. determine the average normal stress at points D and E and represent the stress on a volume element located at each of these points B 6kN 4 kN 8KN
Answer #1

average normal stress at D = Force / Area

= F/pi/4 * (d2^2 - d1^2)

= 4000 / pi/4* (0.028^2 - 0.02^2)

= 1.33*10^7 N/m^2

average normal stress at E = 8000 / pi/4 * 0.012^2

= 7.07*10^7 N/m^2